# GShield: Mitigating Poisoning Attacks in Federated Learning

### v1

@@ 3.1 Federated Learning
Federated Learning (FL, hereafter) is a distributed approach developed to address optimization problems in an environment where a central server coordinates the learning process across a large number of decentralized devices. The devices, known as clients or workers, collaboratively train a Global Model (GM) by processing their local data samples and actively preserving data privacy by avoiding the exchange of raw data. We assume that each client $P_{i}$, for all $i\in[1,N]$, possesses a local training dataset $\mathcal{D}_{i}$. These datasets are mutually disjoint and follow a non-identically independent distributed (non-IID) pattern, i.e., $\mathcal{D}_{i}\cap\mathcal{D}_{j}=\emptyset$ for all $i,j\in[1,N]$ such that $P_{i}\neq P_{j}$. The FL objective seeks to minimize a global loss function $f(w)$, defined as the average of local objectives across $N$ clients:
@@
where $\eta$ is the local learning rate, and $\mathcal{B}_{i,e}$ denotes the mini-batch sampled at epoch $e$ on client $i$. The updated model $w_{i}^{t,E}$ is then sent to the central server for aggregation.
@@ 4.2 Defense Approach Description
We outline our method in Algorithm 1. At the beginning of round $t$, the server randomly selects a set $S^{t}$ of $K$ clients and sends the current global model $w^{t-1}$ to all clients in $S^{t}$. Each selected client $i\in S^{t}$ then performs local training on its dataset $D_{i}$. If client $i$ is an adversary (denoted as $Adv$), it applies a label-flipping attack to its dataset $D_{i}$. After completing the local training, each client computes its gradient $g_{i}^{t}=\nabla\mathcal{L}(w^{t-1};D_{i})$ using stochastic gradient descent and returns this gradient to the server.
@@ 5.4 Performance Analysis of GShield in Tabular Dataset
In this section, we examine the robustness of the proposed GShield method under various safe zones configurations and compare its performance with that of the state-of-the-art method. To assess the sensitivity to the safe period, we configured $SafeRound$ to 5, 10, 15. We used the same experimental setup for all state-of-the-art approaches. Specifically, no malicious participants were introduced until the prescribed $SafeRound$, after which malicious participants were added in subsequent rounds. Table 2 presents the performance of our GShield framework compared to state-of-the-art methods on two tabular benchmark datasets under varying PMR attack ratios and different safe rounds for brevity. GShield consistently outperforms other methods across all datasets and attack scenarios. For CIC-Darkent2020, at lower safe rounds (e.g., 5 rounds), all approaches exhibit relatively low recall values due to the difficulty of detecting malicious updates in a highly non-IID environment. Nevertheless, the proposed GShield demonstrates effectiveness, achieving a 5.4% higher recall than the best baseline (FLAME) at 10% PMR, highlighting its enhanced capability to identify adversarial clients even during early aggregation phases.
@@
For the CSE-CIC-IDS2018 dataset, we observed that GShield achieved approximately $64.884\%$ to $43.253\%$ higher performance for the $SRecall$ than other methods when $SafeRound$ was set to 5 with a PMR of $5\%$. In the case of a $safeRound$ of 5 with a PMR of $25\%$, the $SRecall$ for GShield showed an improvement of approximately $76\%$ compared with other methods, while all other approaches were closer to $0\%$. This result demonstrates GShield’s ability to detect malicious participants effectively while maintaining the $SRecall$. Similarly, in this setting, GShield maintained an $F1Score$ of approximately $87\%$ across all cases, which was higher than the other approaches, whose scores ranged from $70\%$ to $72\%$. A similar trend was observed in other cases as well. The proposed GShield maintained the performance of the global model by effectively detecting both malicious and low-quality models during each round. Under the worst-case attack setting ($SafeRound=15$ and $PMR=15\%$), the $SRecall$ reached $91.213\%$, which is approximately $33\%$ to $41\%$ higher than that of other methods. Meanwhile, the global model achieved an $F1Score$ of $0.9740$.
@@ 5.5 Performance Analysis of GShield in Image Datasets
Table 3 illustrates the performance of the proposed GShield method under different $SafeRound$ and attack scenario settings. In the FashionMNIST dataset, for the $SRecall$ metric, GShield achieved approximately $2–3\%$ improvement over the competing methods for $PMR=5\%$ and $SafeRound=5$. Similarly, we observed that the $F1Score$ of the proposed GShield and FLAME remained almost identical. However, compared to other approaches, GShield improved approximately $2–4\%$. As the poisoning intensity increased, the $SRecall$ of all approaches declined significantly, although most methods attempted to maintain a stable $F1Score$. Notably, GShield still demonstrated an improvement of about $4–12\%$ when the $PMR$ increased to $25\%$, highlighting its robustness in handling higher levels of data poisoning.
Furthermore, as the $SafeRound$ increased, GShield consistently outperformed the other approaches. However, in most cases, FLAME’s performance was comparable to GShield, primarily because FLAME is capable of distinguishing anomalous models even under a high proportion of malicious participants. Nevertheless, GShield outperformed all other methods in both metrics, achieving approximately $1–3\%$ higher performance in $SRecall$, while maintaining a comparable $F1Score$ under the worst-case attack scenario, where other methods exhibited significant degradation.

### v2

@@ 3.1 Federated Learning
Federated Learning (FL, hereafter) is a distributed approach developed to address optimization problems in an environment where a central server coordinates the learning process across a large number of decentralized devices. The devices, known as clients or workers, collaboratively train a Global Model (GM) by processing their local data samples and actively preserving data privacy by avoiding the exchange of raw data. We assume that each client $P_{i}$, for all $i\in[1,N]$, possesses a local training dataset $\mathcal{D}_{i}$. These datasets are mutually disjoint and follow a non-identically independent distributed (non-IID) pattern, i.e., $\mathcal{D}_{i}\cap\mathcal{D}_{j}=\emptyset$ for all $i,j\in[1,N]$ such that $P_{i}\neq P_{j}$. The FL objective seeks to minimize a global loss function $f(w)$, defined as the average of local objectives across NN clients:
@@
where $\eta$ is the local learning rate, and $\mathcal{B}_{i,e}$ denotes the mini-batch sampled at epoch $e$ on client ii. The updated model $w_{i}^{t,E}$ is then sent to the central server for aggregation.
@@ 4.2 Defense Approach Description
We outline our method in Algorithm 1. At the beginning of round $t$, the server randomly selects a set $S^{t}$ of $K$ clients and sends the current global model $w^{t-1}$ to all clients in $S^{t}$. Each selected client $i\in S^{t}$ then performs local training on its dataset $D_{i}$. If client $i$ is an adversary (denoted as AdvAdv), it applies a label-flipping attack to its dataset $D_{i}$. After completing the local training, each client computes its gradient $g_{i}^{t}=\nabla\mathcal{L}(w^{t-1};D_{i})$ using stochastic gradient descent and returns this gradient to the server.
@@ 5.4 Performance Analysis of GShield in Tabular Dataset
In this section, we examine the robustness of the proposed GShield method under various safe zones configurations and compare its performance with that of the state-of-the-art method. To assess the sensitivity to the safe period, we configured SafeRoundSafeRound to 5, 10, 15. We used the same experimental setup for all state-of-the-art approaches. Specifically, no malicious participants were introduced until the prescribed SafeRoundSafeRound, after which malicious participants were added in subsequent rounds. Table 2 presents the performance of our GShield framework compared to state-of-the-art methods on two tabular benchmark datasets under varying PMR attack ratios and different safe rounds for brevity. GShield consistently outperforms other methods across all datasets and attack scenarios. For CIC-Darkent2020, at lower safe rounds (e.g., 5 rounds), all approaches exhibit relatively low recall values due to the difficulty of detecting malicious updates in a highly non-IID environment. Nevertheless, the proposed GShield demonstrates effectiveness, achieving a 5.4% higher recall than the best baseline (FLAME) at 10% PMR, highlighting its enhanced capability to identify adversarial clients even during early aggregation phases.
@@
For the CSE-CIC-IDS2018 dataset, we observed that GShield achieved approximately $64.884\%$ to $43.253\%$ higher performance for the SRecallSRecall than other methods when SafeRoundSafeRound was set to 5 with a PMR of $5\%$. In the case of a safeRoundsafeRound of 5 with a PMR of $25\%$, the SRecallSRecall for GShield showed an improvement of approximately $76\%$ compared with other methods, while all other approaches were closer to $0\%$. This result demonstrates GShield’s ability to detect malicious participants effectively while maintaining the SRecallSRecall. Similarly, in this setting, GShield maintained an $F1Score$ of approximately $87\%$ across all cases, which was higher than the other approaches, whose scores ranged from $70\%$ to $72\%$. A similar trend was observed in other cases as well. The proposed GShield maintained the performance of the global model by effectively detecting both malicious and low-quality models during each round. Under the worst-case attack setting ($SafeRound=15$ and $PMR=15\%$), the SRecallSRecall reached $91.213\%$, which is approximately $33\%$ to $41\%$ higher than that of other methods. Meanwhile, the global model achieved an $F1Score$ of $0.9740$.
@@ 5.5 Performance Analysis of GShield in Image Datasets
Table 3 illustrates the performance of the proposed GShield method under different SafeRoundSafeRound and attack scenario settings. In the FashionMNIST dataset, for the SRecallSRecall metric, GShield achieved approximately $2–3\%$ improvement over the competing methods for $PMR=5\%$ and $SafeRound=5$. Similarly, we observed that the $F1Score$ of the proposed GShield and FLAME remained almost identical. However, compared to other approaches, GShield improved approximately $2–4\%$. As the poisoning intensity increased, the SRecallSRecall of all approaches declined significantly, although most methods attempted to maintain a stable $F1Score$. Notably, GShield still demonstrated an improvement of about $4–12\%$ when the PMRPMR increased to $25\%$, highlighting its robustness in handling higher levels of data poisoning.
Furthermore, as the SafeRoundSafeRound increased, GShield consistently outperformed the other approaches. However, in most cases, FLAME’s performance was comparable to GShield, primarily because FLAME is capable of distinguishing anomalous models even under a high proportion of malicious participants. Nevertheless, GShield outperformed all other methods in both metrics, achieving approximately $1–3\%$ higher performance in SRecallSRecall, while maintaining a comparable $F1Score$ under the worst-case attack scenario, where other methods exhibited significant degradation.